# Online Episodic Memory Visual Query Localization with Egocentric Streaming Object Memory

### v1

@@ A.2.2 Object Discovery (OD) module
Figure 8 reports qualitative examples, reporting the predictions of the discussed models on a frame from the validation set of EgoTracks. Results are consistent with the evaluation metrics in Table 9, Faster R-CNN (A) is the model with the highest recall, indeed it is the model with the highest number of bounding boxes containing both correctly localized objects and errors. YOLO trained on EgoTracks (B) has fewer predictions than Faster R-CNN, but they are more accurate. Training YOLO on EgoTracks (C) shows similar results to Faster R-CNN, with the second densest predictions, translated into the second-highest recall model. YOLO trained on EgoObjects fine-tuned in EgoTracks (D) has the most balanced results; predictions are well localized, reflecting the best performing model in the HM metric. GroundingDino (E) contains several correctly regressed objects, but as shown in the frame, no ground truth annotations were found. Hand-Object Detector (F) has the worst performance, containing only two predictions, one of which is related to the interacted object. The discovery mechanism of this detector leads to the lowest values for both precision and recall. The predictions of VideoAgent (G) are partially correct, despite the presence of a few wrong bounding boxes. Finally the predictions marked with H are the ground truth ones.
@@
Given these results, in the following experiments, we consider the YOLO detector trained on EgoObjects and EgoTracks as a representative of trained methods and both GroundingDino and Hand-Object-Detector as representative of off-the-shelf methods. These are referred to as ODyolo and ODgdino respectively in the main paper.
@@
YOLO EgoTracks 0.1983 0.1634 0.1791
YOLO EgoObjects 0.0247 0.3780 0.0463
YOLO EgoObjects + EgoTracks∗∗ 0.1358 0.3067 0.1882
@@
Figure 8: Visualization of predictions from the object detection models. Here we qualitatively show the performance of the different object detectors over the same sampled frame. A: Faster R-CNN on EgoTracks, B: YOLO on EgoTracks, C: YOLO on EgoObjects, D: YOLO on EgoObjects, fine-tuned in EgoTracks, E: GroundingDino, F: is the Hand-Object Detector G: RT-DETR, H: Ground-truth annotations.
@@ A.3 Query Retrieval and Localization (QRL)
In both cases, for each object i, the similarity scores are averaged into a single score $r_{i}$. If the maximum score $r_{i}$ exceeds the threshold $\lambda_{ref}$ = 0.5, the object $\mathbf{O}_{i}$ is considered a match for query $\mathbf{Q}$. Once match is identified, the most recent sequence of contiguous bounding-boxes from $\mathbf{O}_{i}$’s spatio-temporal history is retrieved and returned as the response track r.
@@ Appendix B Online Object Memory Baselines
Here we provide details of the implementations of the baselines used for comparison in Table 2.
@@ B.1 AMEGO
The AMEGO framework [17], was employed to construct an episodic memory model, serving as a performance benchmark for our proposed approach. We utilized ce683bd version of the AMEGO codebase, available at  [^6] https://github.com/gabrielegoletto/AMEGO.git, with minor modifications to accommodate the EGO4D dataset structure. Our implementation focused specifically on the Hand-Object Interaction (HOI) module of AMEGO. The Location Segmentation module was excluded from our analysis as it fell outside the scope of our research objectives. To ensure a fair and consistent comparison, we maintained the original hyperparameters and thresholds to align with the baseline model published on the official GitHub repository. Since their experiment has been made using a different dataset, our modifications were limited to data input/output functions to handle the EGO4D data format, without altering the core algorithmic components.
@@ B.2 VideoAgent
The VideoAgent framework [11], has been utilized as a comparative baseline for our episodic memory model. We employed the 21e4eb7 version of the VideoAgent repository [^7] https://github.com/YueFan1014/VideoAgent.git. Our implementation focused on the Object Memory module, and we opted to omit the captioning module since it was not pertinent to our research scope. To maintain experimental integrity, we preserved the original configuration parameters.
@@ B.3 Memory Representation
As with the other OMP configurations discussed in this paper, the object tracklets produced by VideoAgent and Amego OMP are stored in memory using the MR1* representation, as it provides the best balance between accuracy and memory efficiency (see Table 6). Specifically, for each $\mathbf{O}_{i}$ found by either AMEGO or VideoAgent object detection and tracking algorithms, $\mathcal{M}_{ego}$ retains the RGB patch ($\phi_{i,t}$) and the $\mathbf{b}_{i,t}$ at the time of object discovery, along with two RGB patches from the most recent response track. These patches correspond to the first and last valid patches identified by the classifier within the most recent sequence of bounding-boxes. For all other temporal locations not covered above, $\mathcal{M}_{ego}$ stores only the spatio-temporal information provided by $\mathbf{b}_{i,t}$.
@@ Appendix C Evaluation Dataset
Due to the high storage requirements for the validation split of the Ego4D VQ2D benchmark [18] and the significant time needed for experiments on the full dataset, we opted to use a subset of the validation clips for our evaluations. Specifically, we randomly selected 115 video clips uniformly, containing 450 visual queries along with their corresponding ground-truth response tracks and tracklets. This subset consisted of a total of 208365 frames. To maintain consistency with the original frame rate of the EGO4D dataset [18], we ran the OMP algorithm at 5 frames per second.
@@ Appendix D Additional Results
OMP Configuration tAP25 $\uparrow$ stAP25 $\uparrow$ Succ $\uparrow$ Size $\downarrow$
OD OT
ODoracle OTegostark 0.18 0.05 31.91 40.3 MB
ODyolo OTegostark 0.0003 0.002 4.02 209.3 MB
ODoracle+yolo OTegostark 0.10 0.02 23.67 229.4 MB
@@
Figure 9: Additional qualitative examples. Successful retrievals of three different queries $\mathbf{Q}$ from $\mathcal{M}_{ego}$ are shown using (a) the oracle OMP algorithm (ODoracle - OToracle) and (b) the real-world OMP (ODyolo - OTegostark). The figure highlights the visual representations stored in $\mathcal{M}_{ego}$ and matched during retrieval, their respective timestamps, the duration of the response track $\mathbf{r}$ returned by QRL, and the extent of the ground-truth response track (the rectangle with dashed outline).

### v2

@@ A.2.2 Object Discovery (OD) module
Figure 8 reports qualitative examples, reporting the predictions of the discussed models on a frame from the validation set of EgoTracks. Results are consistent with the evaluation metrics in Table 9, Faster R-CNN (A) is the model with the second-highest recall, indeed it is the model with the highest number of bounding boxes containing both correctly localized objects and errors. YOLOv10 trained on EgoTracks (B) has fewer predictions than Faster R-CNN, but they are more accurate. Training YOLOv10 on EgoTracks (C) shows similar results to Faster R-CNN, with the second-densest predictions, translated into the second-highest recall model. YOLOv10 trained on EgoObjects fine-tuned in EgoTracks (D) has the most balanced results; predictions are well localized, reflecting the best-performing model in the HM metric. YOLOWorld (E) achieves results comparable to Faster R-CNN (A), demonstrating the highest recall with dense object predictions and minimal errors, which can be further reduced by increasing the threshold. iGroundingDino (F) contains several correctly regressed objects, but as shown in the frame, no ground truth annotations were found. Hand-Object Detector (G) has the worst performance, containing only two predictions, one of which is related to the interacted object. The discovery mechanism of this detector leads to the lowest values for both precision and recall. The predictions of VideoAgent (H) are partially correct, despite the presence of a few wrong bounding boxes. Finally the predictions marked with I are the ground truth ones.
@@
Given these results, in the following experiments, we consider the YOLOv10 detector trained on EgoObjects and EgoTracks as a representative of trained methods and both GroundingDino and Hand-Object-Detector as representative of off-the-shelf methods. These are referred to as ODyolo and ODgdino respectively in the main paper.
ODoracle
. The oracular object detector is obtained using the ground truth object annotations in EgoTracks as valid detections. As previously mentioned, EgoTracks is aligned to the Episodic Memory benchmark of Ego4D, so EgoTracks detections include valid queries.
@@
YOLOv10 EgoTracks 0.1983 0.1634 0.1791
YOLOv10 EgoObjects 0.0247 0.3780 0.0463
YOLOv10 EgoObjects + EgoTracks∗∗ 0.1358 0.3067 0.1882
YOLOWorld - 0.0207 0.5969 0.0401
@@
Figure 8: Visualization of predictions from the object detection models. Here we qualitatively show the performance of the different object detectors over the same sampled frame. A: Faster R-CNN on EgoTracks, B: YOLOv10 on EgoTracks, C: YOLOv10 on EgoObjects, D: YOLOv10 on EgoObjects, fine-tuned in EgoTracks, E: YOLOWorld, F: GroundingDino, G: is the Hand-Object Detector H: RT-DETR, I: Ground-truth annotations.
@@ A.3 Query Retrieval and Localization (QRL)
In both cases, for each object i, the similarity scores are averaged into a single score $r_{i}$. If the maximum score $r_{i}$ exceeds the threshold $\lambda_{ref}$ = 0.5, the object $\mathbf{O}_{i}$ is considered a match for query $\mathbf{Q}$. Once a match is identified, the most recent sequence of contiguous bounding boxes from $\mathbf{O}_{i}$’s spatio-temporal history is retrieved and returned as the response track r.
A.4 ESOM Memory Overhead
The GPU memory overhead is 423MB for ODyolo and 2,350MB for OTegostark when tracking $10$ objects concurrently, 2,320MB for OTmasa, and 2,474MB for QRLsiamrcnn.
@@ Appendix B Online Object Memory Baselines
Here we provide details of the implementations of the baselines used for comparison.
@@ B.1 AMEGO
The AMEGO framework [21], was employed to construct an episodic memory model, serving as a performance benchmark for our proposed approach. We utilized ce683bd version of the AMEGO codebase, available at https://github.com/gabrielegoletto/AMEGO.git, with minor modifications to accommodate the EGO4D dataset structure. Our implementation focused specifically on the Hand-Object Interaction (HOI) module of AMEGO. The Location Segmentation module was excluded from our analysis as it fell outside the scope of our research objectives. To ensure a fair and consistent comparison, we maintained the original hyperparameters and thresholds to align with the baseline model published on the official GitHub repository. Since their experiment has been made using a different dataset, our modifications were limited to data input/output functions to handle the EGO4D data format, without altering the core algorithmic components.
@@ B.2 VideoAgent
The VideoAgent framework [15], has been utilized as a comparative baseline for our episodic memory model. We employed the 21e4eb7 version of the VideoAgent repositoryhttps://github.com/YueFan1014/VideoAgent.git. Our implementation focused on the Object Memory module, and we opted to omit the captioning module since it was not pertinent to our research scope. To maintain experimental integrity, we preserved the original configuration parameters.
@@ B.3 Memory Representation
As with the other OMP configurations discussed in this paper, the object tracklets produced by VideoAgent and Amego OMP are stored in memory using the MR1* representation, as it provides the best balance between accuracy and memory efficiency (see Table 6). Specifically, for each $\mathbf{O}_{i}$ found by either AMEGO or VideoAgent object detection and tracking algorithms, $\mathcal{M}_{ego}$ retains the RGB patch ($\phi_{i,t}$) and the $\mathbf{b}_{i,t}$ at the time of object discovery, along with two RGB patches from the most recent response track. These patches correspond to the first and last valid patches identified by the classifier within the most recent sequence of bounding boxes. For all other temporal locations not covered above, $\mathcal{M}_{ego}$ stores only the spatio-temporal information provided by $\mathbf{b}_{i,t}$.
@@ Appendix C Evaluation Dataset
Due to the high storage requirements for the validation split of the Ego4D VQ2D benchmark [22] and the significant time needed for experiments on the full dataset, we opted to use a subset of the validation clips for our evaluations. Specifically, we randomly selected 115 video clips uniformly, containing 450 visual queries along with their corresponding ground-truth response tracks and tracklets. This subset consisted of a total of 208,365 frames. To maintain consistency with the original frame rate of the EGO4D dataset [22], we ran the OMP algorithm at 5 frames per second.
OMP Configuration %tAP25 $\uparrow$ %stAP25 $\uparrow$ %Succ $\uparrow$ Size $\downarrow$
OD OT
ODoracle OTegostark 18.3 5.1 31.91 591.4 MB
ODyolo OTegostark 0.03 0.2 4.02 1.7 GB MB
ODoracle+yolo OTegostark 10.6 2.3 23.67 2.4 GB
@@ Appendix D Additional Results
Random performance The performance of ESOM has been evaluated by implementing a QRL method that randomly (w. uniform dist.) selects a response track from $\mathcal{M}_{ego}$. This achieves a Succ. score of 27.29 (ODoracle+OToracle) and of 0.44 (ODyolo+OTegostark), while ESOM achieves 81.92 and 4.02 respectively. ESOM performance, even in suboptimal OMP configuration (ODyolo+OTegostark) are still better than random selection. This indicates that the method effectively leverages available information to enhance performance.
Constant memory budget We explored a “constant-budget” version of ESOM, in which older and low-confidence objects are pruned as soon as a predefined memory limit is reached. Results (Figure 9) demonstrate that ESOM isn’t particularly affected by memory constraints.
Figure 9: Impact of memory-budget on ESOM performances. Success rate and query time for two ESOM variants under varying memory budgets (250–1000 MB). Results show that performance remains stable as older or low-confidence objects are pruned to meet memory constraints, indicating ESOM’s robustness under limited memory conditions.
@@
Figure 10: Additional qualitative examples. Successful retrieval of thre different queries $\mathbf{Q}$ from $\mathcal{M}_{ego}$ using the oracle OMP algorithm (ODoracle - OToracle) and the real-world OMP (ODyolo - OTegostark). Sample frames from each track are shown with corresponding bounding boxes whose patches $\phi_{i,t}$ were matched during retrieval. The frames in gray correspond to the moment of the object’s discovery. The green bar visualizes the temporal extent of the ground-truth track, with timestamps indicating its duration.